# The Dynamical History of the Kepler-221 Planet System

### v1

@@ 5.2 Orbital Expansion: specific scenarios
In this subsection, we conduct parameter studies for the two mechanisms outlined above: (i) outward migration of the b/c/e system; or (ii) a change in the masses of the planets.
@@ 5.2.1 Outward-moving planet b
Figure 7: Time dependence of torque on b for different physical mechanisms. Dashed lines: minimum dynamical torque on (blue) or mass loss rate for (green) planet b required to induce a high enough migration of planet c, which ensures sustained orbital expansion of the b/c/e system. These torques decay with time according to the models by Wang & Lin (2023) (mass loss) and Ahuir et al. (2021) (dynamical torques). For reference, the green solid line represents the torque on b due to a 2% mass loss (Wang & Lin, 2023) and the blue solid line fits the dynamical tide model of Ahuir et al. (2021). The black line represents the torque due to equilibrium tides assuming $Q_{\star}^{\prime}=10^{6}$.
@@
$\Gamma_{\mathrm{ML}}=\frac{L_{\mathrm{b}}}{\tau_{\mathrm{m}}}\frac{A\exp(-t/\tau_{\mathrm{m}})}{1+A\exp(-t/\tau_{\mathrm{m}})}\approx\Gamma_{\mathrm{ML,0}}\exp(-t/\tau_{\mathrm{m}})$ (12)
@@
$\Gamma_{\mathrm{DT}}=\Gamma_{\mathrm{DT,0}}\left(\frac{t+\tau_{\mathrm{d}}}{\tau_{\mathrm{d}}}\right)^{-2}$ (13)
where $\Gamma_{\mathrm{DT,0}}$ is the initial dynamical torque and $\tau_{\mathrm{d}}=t_{\Gamma}$ is the time after which the dynamical tides rapidly decay. For solar-likes star like Kepler-221 we put $\tau_{\mathrm{d}}=460\,\mathrm{Myr}$ (Ahuir et al., 2021). Figure 7 plots these expressions for $\Gamma_{\mathrm{ML}}$ and $\Gamma_{\mathrm{DT}}$ by the solid green and blue lines, respectively. The torque due to the dynamical tide on planet b is represented by the blue solid line (Ahuir et al., 2021), which is much stronger than the equilibrium tides represented by the black line in Figure 7 (assuming $Q_{\star}^{\prime}=10^{6}$).
@@ 5.2.2 Mass Constraint on b, c, and e
A more direct way to avoid crossing the c/d/e (3,5,8) 3BR is to assume different masses of the planets. The masses of planets b, c, and e determine the slope of expansion of $P_{\mathrm{d}}/P_{\mathrm{c}}$ and $P_{\mathrm{e}}/P_{\mathrm{d}}$ defined as $a_{\mathrm{cde}}$ in Equation (10). To reach the observed period ratio at the end of the orbital expansion phase, the slope should satisfy $a_{\mathrm{cde}}>1.56$ (see Section 5 and Appendix C). Therefore, if there are no additional mechanisms (e.g. Section 5.2.1) acting on the planets, we can constrain the masses and densities of the planets with the current period ratio in the orbital expansion phase.
@@
Figure 8 shows the constraints on the mass and density ratios in the orbital expansion phase in which the migration of planet d in the orbital expansion phase has been neglected. Successful orbital expansion is only possible if the combination of mass or density ratios falls in the green region, which corresponds to $a_{\mathrm{cde}}>1.56$. In general, this requires a more massive planet c than planet e (e.g., mass model M2a or M2d in Table 2, green dot in Figure 8). Because mass model M2a and M2d have the same mass for planets b, c, and e, only one dot is presented in Figure 8. Initial masses according to the mass-radius relationship in Chen & Kipping (2017) would result in too small $a_{\mathrm{cde}}$ for successful orbital expansion (M1 in Table 2, blue dot in Figure 8) and require additional mechanisms (e.g., stronger damping on planet d and outward torque on planet b, see Section 5).
@@ 6 Discussion
I 4.1 1st-order resonance chain with five planets large scale migration $\bullet$ formation of 1st-order 2BR is a natural result of convergent migration
@@
III 4.3 3BR reformation 1. no effects -- 0th-order b/c/e 3BR never reforms
@@
IV 5.2 avoid c/d/e 3BR 1. positive torque on planet b $\bullet$ mechanism can also be applied to Phase III
@@ 6.1 Assessment of the Model
In this work, we have investigated a number of mechanisms to explain the observational constraints that the Kepler-221 system presents us. A summary of the comments for each of these mechanisms is listed in Table 4 in chronological order. In Phase I and II, the merging of planets $\mathrm{d}_{1}$ and $\mathrm{d}_{2}$ is a natural result of the 1st-order resonance breaking after disk dispersal, and a quick merging within hundreds of years is achieved with high eccentricities for planets after disk dispersal (see Section 4.1).
Perhaps the most critical phase for the collision scenario is the re-establishment of the 6:3:1 nonadjacent 3BR between planets b, c, and e. We found that reformation is not spontaneous (Petit, 2021), but instead requires either stronger damping on planet c or a positive torque on planet b. The latter scenario is favored because the outward migration of planet b also promotes a successful subsequent orbital expansion phase, which ensures that planets c, d, and e can avoid 3BR encounters during Phase IV (see Section 5.2.1). Still, resonance reformation is not guaranteed, but stochastic, and successful only in ${\sim}17\%$ of our simulations (see Section 4.3).
It is likely that planets b, c, and e have experienced significant orbital expansion while locked in the 3BR configuration. If the system is young, as suggested by Berger et al. (2018), eccentricity tides alone are unlikely to be the sole driver responsible for the 3BR expansion. Therefore, additional damping mechanisms, such as obliquity tides (Goldberg & Batygin, 2021), are required to accelerate the expansion. However, a too-strong dissipation would hinder the reformation of b/c/e 3BR in Phase III. This tension between Phase III (which favors weak damping) and Phase IV (which favors robust damping) is discussed below in Section 6.3.
@@ 6.2 Mass constraints
Our modeling provides constraints on the yet undetermined masses of the Kepler-221 planets (Berger et al., 2018). The first clue comes from the formation scenario (Phase II-III), where it was postulated that planet d arises from the merger of two planets. The merging could have attributed to significant atmosphere loss but little mass loss (Ghosh et al., 2024; Dou et al., 2024). If so, it is expected that the density of planet d is higher, which would make planet d the most massive planet in the current system. The reformation of 3BR in the post-collision phase is still viable with planet d twice as massive as other planets (see Section 4.4). Furthermore, once the b/c/e resonance is established, the long-term evolution of the system (Phase IV) is mostly independent of planet d’s mass (see Appendix D). Therefore, a more massive planet d remains a possibility.
@@
The second stronger clue comes from the orbital expansion phase. This clue is stronger as it is independent of the formation model. In the orbital expansion phase, a slope of $a_{\mathrm{cde}}>1.56$ in the c/d/e orbital plane (see Section 5.1) is required, which constrains the masses and densities of planets b, c, and e in the absence of additional dissipative forces or torques on planets (see Section 5.2.2). From this, similar masses with a slightly more massive planet c are expected, such as in model M2a, consistent with the peas-in-a-pod model (Weiss & Petigura, 2020; Weiss et al., 2023). This model renders the densities of the planets progressively decrease with distance, $\rho_{b}>\rho_{c}>\rho_{e}$ (see Figure 8). A possible origin of the similar masses and the decreasing density of the planets in the system could be due to photoevaporation. Inflated low-mass planets with orbits of a few tenths of an AU could experience a boil-off phase after disk dispersal, in which the planets lose their envelope and contract quickly with only a few of their original envelope left after the disk dispersal (Owen & Wu, 2016). Afterwards, an X-ray and EUV-dominated photoevaporation could further erode the planet atmosphere with a timescale of approximately $100\,\mathrm{Myr}$ (Owen & Wu, 2013; Lopez & Fortney, 2013). One example of systems with photoevaporation-dominated planets is Kepler-36, which hosts two close-in planets (closer than $0.15\mathrm{au}$) with similar atmosphere fractions when they are born and the inner planet has lost almost all of its atmosphere (Carter et al., 2012; Lopez & Fortney, 2013; Owen, 2019). A similar photoevaporation process could have happened in the Kepler-221 system for planet b, resulting in a planet b of terrestrial density and puffier outer planets. An additional advantage of photoevaporation on planet b is that the associated rapid mass loss would contribute to the outward migration of planets b and c, which facilitates success in Phase III and Phase IV of our model (see Sects. 5.2.1 and 4.3).
@@
Figure 9: Mass-radius diagram of the Kepler-221 planets with mass models listed in Table 2 along with the exoplanets sample. Black dots and error bars represent the masses of planets according to mass model M2a of Table 2 and radii according to Table 1. The blue line shows the masses of the planets according to mass model M2 of Table 2. The gray lines provide the mass-radius relation according to Chen & Kipping (2017). The green line indicates the mass of d if it would be twice as massive as the other planets, following mass model M2d of Table 2. The background dots and error bars show the masses and radii of other exoplanets, with those in red representing planets in a 1st-order resonance chain involving at least three planets. The gray area shows the radius valley between $1.5R_{\oplus}$ and $2.0R_{\oplus}$ according to Fulton et al. (2017).
@@
Assuming model M2a of Table 2, Figure 9 plots the planets in the Kepler-221 system along with the planets of the exoplanet sample for which masses and radius are available [^5] https://exoplanetarchive.ipac.caltech.edu. The plot shows that assuming similar masses following the peas-in-a-pod model (Weiss & Petigura, 2020; Weiss et al., 2023) in M2a in Table 2 results in planet c and e puffier than Neptune. However, such an assumption is not unreasonable because similar super-puff planets exist in the exoplanet sample, especially in resonance chains (red background dots in Figure 9). For example, TOI-178 hosts six planets in a 1st-order resonance except for the innermost planets (Leleu et al., 2021b). Among these planets, TOI-178 d and g are super-puff planets with radii similar to and density lower than the planets c and e estimated in M2a in Table 2 (Leleu et al., 2021b; Delrez et al., 2023). Kepler-51 hosts three planets in a 3:2:1 resonance chain, and two of the planets are super-puff planets with masses similar to and radii much larger than the Kepler-221 estimation for planets c and e in mass models M2a. Therefore, the outer planets in the Kepler-221 system could potentially be another sample of super-puff planets in resonance.
@@
It is worthwhile to conduct RV measurements on the Kepler-221 system with facilities like Keck (Vogt et al., 1994). The significance of the RV detection is as follows: (i) if planet d turns out to be more massive than the other planets, it suggests a collisional origin. (ii) If the mass ratios of the planets b, c, and e fall in the green region in Figure 8, it offers the most natural scenario to explain the sustained orbital expansion phase (see Section 5). And (iii) it can provide evidence that planets c and e belong to the super-puff planets, while planet b is on the other side of the radius valley (Fulton et al., 2017) (see Figure 9).
@@ 6.3 Age constraint
We define the time required to expand the b, c, and e 3BR from the integer period ratio to the observed period ratio as $t_{\mathrm{exp}}$. In our simulation, $t_{\mathrm{exp}}$ mainly depends on the tidal damping strength on the planets (determined by $Q_{\mathrm{phy}}$) and is hardly affected by other parameters like the mass ratios of the planets. For example, panels e and f in Figure 6 show a successful expansion with masses of the planets according to mass model M2a in Table 2 and the same $Q_{\mathrm{phy}}$ for all planets. From the orbital expansion phase simulations (see Section 5), the relation between $t_{\mathrm{exp}}$ and $Q_{\mathrm{phy}}$ can be expressed as:
$t_{\mathrm{exp}}\approx 380Q_{\mathrm{phy}}\,\mathrm{Myr}$ (14)
@@
Such results conflict the the formation model of Kepler-221 in two aspects. Firstly, The value of $Q_{\mathrm{phy}}$ is over two orders of magnitudes smaller than the typical value of the tidal quality factor $Q$ for terrestrial planets (Lee et al., 2013; Silburt & Rein, 2015), indicating the existence of other much more efficient dissipation mechanisms to speed up the simulation. Secondly, the reformation process of 3BR requires gentle convergent migration between planets b/c/e corresponding to relatively large $Q_{\mathrm{phy}}$ (in Phase III we apply $Q_{\mathrm{phy}}>3$, see Table 3 and Section 4.4). This conflicts with the $Q_{\mathrm{phy}}<2$ constraint derived from the age of the system.
@@
Table 6: Dynamical configuration of Kepler-402 planetary system according to Rowe et al. (2014).
@@
Goldberg & Batygin (2021) point out that the timescale puzzle could be solved with some additional dissipation mechanism is required (e.g., obliquity tides; Millholland & Laughlin 2019). Obliquity tides occur when planets have a large axial tilt (obliquity) (Millholland & Laughlin, 2019) and have been invoked to speed up the orbital expansion in the Kepler-221 system (Goldberg & Batygin, 2021). A non-zero obliquity for planets can be maintained if the planet is in Cassini state where it reaches a secular spin-orbit resonance with synchronized precession of planetary spin and orbital angular momentum. The total energy dissipation rate for a planet with non-zero eccentricity and obliquity can be expressed as (Millholland & Laughlin, 2019):
$\dot{E}(e,\epsilon)=\frac{2K}{1+\cos^{2}\epsilon}\left[\sin^{2}\epsilon+e^{2}\left(7+16\sin^{2}\epsilon\right)\right]$ (15)
@@
Planets are likely to have low obliquity during the disk phase as their angular momentum and spin vectors are aligned with the disk. But during the brief phase of dynamical instability and the resonance reformation afterward, planet b or c could have acquired a mutual inclination with respect to the other planets, which is a requirement for trapping in a Cassini state. Additionally, planet c dominate the tidal expansion process at the same level as planet b due to the large radius of c compared to b (see Equation (6)). Equation (15) demonstrates that even modest obliquity (for example, on planet c, consistent with Section 4.3) can dominate the energy dissipation, i.e., to drive the tidal expansion at rates much higher than what can be obtained from tidal damping alone.

### v2

@@ 5.2 Orbital expansion: Specific scenarios
In this subsection, we conduct parameter studies for the two mechanisms outlined above: (i) outward migration of the (b, c, e) system; or (ii) a change in the masses of the planets.
@@ 5.2.1 Outward-moving planet b
Figure 7: Time dependence of torque on b for different physical mechanisms. The dashed lines represent the minimum dynamical torque on (blue) or mass loss rate for (green) planet b required to induce a high enough migration of planet c, which ensures sustained orbital expansion of the (b, c, e) system. These torques decay with time according to the models by Wang & Lin (2023) (mass loss) and Ahuir et al. (2021) (dynamical torques). For reference, the green solid line represents the torque on b due to a 2% mass loss (Wang & Lin, 2023) and the blue solid line fits the dynamical tide model of Ahuir et al. (2021). The black line represents the torque due to equilibrium tides assuming $Q_{\star}^{\prime}=10^{6}$.
@@
$\Gamma_{\mathrm{ML}}=\frac{L_{\mathrm{b}}}{\tau_{\mathrm{m}}}\frac{A\exp(-t/\tau_{\mathrm{m}})}{1+A\exp(-t/\tau_{\mathrm{m}})}\approx\Gamma_{\mathrm{ML,0}}\exp(-t/\tau_{\mathrm{m}}),$ (12)
@@
$\Gamma_{\mathrm{DT}}=\Gamma_{\mathrm{DT,0}}\left(\frac{t+\tau_{\mathrm{d}}}{\tau_{\mathrm{d}}}\right)^{-2},$ (13)
where $\Gamma_{\mathrm{DT,0}}$ is the initial dynamical torque and $\tau_{\mathrm{d}}=t_{\Gamma}$ is the time after which the dynamical tides rapidly decay. For solar-likes star such as Kepler-221 we put $\tau_{\mathrm{d}}=460\,\mathrm{Myr}$ (Ahuir et al., 2021). Figure 7 plots these expressions for $\Gamma_{\mathrm{ML}}$ and $\Gamma_{\mathrm{DT}}$ by the solid green and blue lines, respectively. The torque due to the dynamical tide on planet b is represented by the blue solid line (Ahuir et al., 2021), which is much stronger than the equilibrium tides represented by the black line in Figure 7 (assuming $Q_{\star}^{\prime}=10^{6}$).
@@ 5.2.2 Mass constraint on b, c, and e
A more direct way to avoid crossing the (c, d, e) (3,5,8) 3BR is to assume different masses of the planets. The masses of planets b, c, and e determine the slope of expansion of $P_{\mathrm{d}}/P_{\mathrm{c}}$ and $P_{\mathrm{e}}/P_{\mathrm{d}}$ defined as $a_{\mathrm{cde}}$ in Equation (10). To reach the observed period ratio at the end of the orbital expansion phase, the slope should satisfy $a_{\mathrm{cde}}>1.56$ (see Section 5 and Appendix C). Therefore, if there are no additional mechanisms (for instance Section 5.2.1) acting on the planets, we can constrain the masses and densities of the planets with the current period ratio in the orbital expansion phase.
@@
Figure 8 shows the constraints on the mass and density ratios in the orbital expansion phase in which the migration of planet d in the orbital expansion phase has been neglected. Successful orbital expansion is only possible if the combination of mass or density ratios falls in the green region, which corresponds to $a_{\mathrm{cde}}>1.56$. In general, this requires a more massive planet c than planet e (for instance, mass model M2a or M2d in Table 2, green dot in Figure 8). Because mass model M2a and M2d have the same mass for planets b, c, and e, only one dot is presented in Figure 8. Initial masses according to the mass-radius relationship in Chen & Kipping (2017) would result in too small $a_{\mathrm{cde}}$ for successful orbital expansion (M1 in Table 2, blue dot in Figure 8) and require additional mechanisms (for instance, stronger damping on planet d and outward torque on planet b, see Section 5).
@@ 6 Discussion
I 4.1 First-order resonance chain with five planets Large scale migration $\bullet$ Formation of first-order 2BR is a natural result of convergent migration
@@
III 4.3 3BR reformation 1. No effects -- Zeroth-order (b, c, e) 3BR never reforms
@@ 3. Positive torque on planet b + Mechanism can also be applied to Phase IV
IV 5.2 Avoid (c, d, e) 3BR 1. Positive torque on planet b $\bullet$ Mechanism can also be applied to Phase III
@@ 6.1 Assessment of the model
In this work, we have investigated several mechanisms to explain the observational constraints that the Kepler-221 system presents to us. A summary of the comments for each of these mechanisms is listed in Table 4 in chronological order. In Phase I and II, the merging of planets $\mathrm{d}_{1}$ and $\mathrm{d}_{2}$ is a natural result of the first-order resonance breaking after disk dispersal, and a quick merging within hundreds of years is achieved with high eccentricities for planets after disk dispersal (see Section 4.1).
Perhaps the most critical phase for the collision scenario is the re-establishment of the 6:3:1 nonadjacent 3BR between planets b, c, and e. We found that reformation is not spontaneous (Petit, 2021), but instead requires either a stronger damping on planet c or a positive torque on planet b. The latter scenario is favored because the outward migration of planet b also promotes a successful subsequent orbital expansion phase, which ensures that planets c, d, and e can avoid 3BR encounters during Phase IV (see Section 5.2.1). Still, resonance reformation is not guaranteed, but stochastic, and successful only in ${\sim}17\%$ of our simulations (see Section 4.3).
It is likely that planets b, c, and e have experienced significant orbital expansion while locked in the 3BR configuration. If the system is young, as suggested by Berger et al. (2018), eccentricity tides alone are unlikely to be the sole driver responsible for the 3BR expansion. Therefore, additional damping mechanisms, such as obliquity tides (Goldberg & Batygin, 2021), are required to accelerate the expansion. However, a too-strong dissipation would hinder the reformation of (b, c, e) 3BR in Phase III. This tension between Phase III (which favors weak damping) and Phase IV (which favors robust damping) is discussed below in Section 6.3.
@@ 6.2 Mass constraints
Our modeling provides constraints on the yet undetermined masses of the Kepler-221 planets (Berger et al., 2018). The first clue comes from the formation scenario (Phase II-III), where it was postulated that planet d arises from the merger of two planets. The merging could have attributed to significant atmosphere loss but little mass loss (Ghosh et al., 2024; Dou et al., 2024). If so, it is expected that the density of planet d is higher, which would make planet d the most massive planet in the current system. The reformation of 3BR in the post-collision phase is still viable with planet d twice as massive as other planets (see Section 4.4). Furthermore, once the (b, c, e) resonance is established, the long-term evolution of the system (Phase IV) is mostly independent of planet d’s mass (see Appendix D). Therefore, a more massive planet d remains a possibility.
@@
The second stronger clue comes from the orbital expansion phase. This clue is stronger as it is independent of the formation model. In the orbital expansion phase, a slope of $a_{\mathrm{cde}}>1.56$ in the (c, d, e) orbital plane (see Section 5.1) is required, which constrains the masses and densities of planets b, c, and e in the absence of additional dissipative forces or torques on planets (see Section 5.2.2). From this, similar masses with a slightly more massive planet c are expected, such as in model M2a, consistent with the peas-in-a-pod model (Weiss & Petigura, 2020; Weiss et al., 2023). This model renders the densities of the planets progressively decrease with distance, $\rho_{b}>\rho_{c}>\rho_{e}$ (see Figure 8). A possible origin of the similar masses and the decreasing density of the planets in the system could be due to photoevaporation. Inflated low-mass planets with orbits of a few tenths of an AU could experience a boil-off phase after disk dispersal, in which the planets lose their envelope and contract quickly with only a few of their original envelope left after the disk dispersal (Owen & Wu, 2016). Afterwards, an X-ray and EUV-dominated photoevaporation could further erode the planet atmosphere with a timescale of approximately $100\,\mathrm{Myr}$ (Owen & Wu, 2013; Lopez & Fortney, 2013). One example of systems with photoevaporation-dominated planets is Kepler-36, which hosts two close-in planets (closer than $0.15\mathrm{au}$) with similar atmosphere fractions when they are born and the inner planet has lost almost all of its atmosphere (Carter et al., 2012; Lopez & Fortney, 2013; Owen, 2019). A similar photoevaporation process could have happened in the Kepler-221 system for planet b, resulting in a planet b of terrestrial density and puffier outer planets. An additional advantage of photoevaporation on planet b is that the associated rapid mass loss would contribute to the outward migration of planets b and c, which facilitates success in Phase III and Phase IV of our model (see Sects. 5.2.1 and 4.3).
@@
Figure 9: Mass-radius diagram of the Kepler-221 planets with mass models listed in Table 2 along with the exoplanets sample. Black dots and error bars represent the masses of planets according to mass model M2a of Table 2 and radii according to Table 1. The blue line shows the masses of the planets according to mass model M2 of Table 2. The gray lines provide the mass-radius relation according to Chen & Kipping (2017). The green line indicates the mass of d if it would be twice as massive as the other planets, following mass model M2d of Table 2. The background dots and error bars show the masses and radii of other exoplanets, with those in red representing planets in a first-order resonance chain involving at least three planets. The gray area shows the radius valley between $1.5R_{\oplus}$ and $2.0R_{\oplus}$ according to Fulton et al. (2017).
@@
Assuming model M2a of Table 2, Figure 9 plots the planets in the Kepler-221 system along with the planets of the exoplanet sample for which masses and radius are available. [^5] https://exoplanetarchive.ipac.caltech.edu The plot shows that assuming similar masses following the peas-in-a-pod model (Weiss & Petigura, 2020; Weiss et al., 2023) in M2a in Table 2 results in planet c and e puffier than Neptune. However, such an assumption is not unreasonable because similar super-puff planets exist in the exoplanet sample, especially in resonance chains (red background dots in Figure 9). For example, TOI-178 hosts six planets in a first-order resonance except for the innermost planets (Leleu et al., 2021b). Among these planets, TOI-178 d and g are super-puff planets with radii similar to and density lower than the planets c and e estimated in M2a in Table 2 (Leleu et al., 2021b; Delrez et al., 2023). Kepler-51 hosts three planets in a 3:2:1 resonance chain, and two of the planets are super-puff planets with masses similar to and radii much larger than the Kepler-221 estimation for planets c and e in mass models M2a. Therefore, the outer planets in the Kepler-221 system could potentially be another sample of super-puff planets in resonance.
@@
It is worthwhile to conduct RV measurements on the Kepler-221 system with facilities such as Keck (Vogt et al., 1994). The significance of the RV detection is as follows: (i) if planet d turns out to be more massive than the other planets, it suggests a collisional origin. (ii) If the mass ratios of the planets b, c, and e fall in the green region in Figure 8, it offers the most natural scenario to explain the sustained orbital expansion phase (see Section 5). And (iii) it can provide evidence that planets c and e belong to the super-puff planets, while planet b is on the other side of the radius valley (Fulton et al., 2017) (see Figure 9).
@@ 6.3 Age constraint
We define the time required to expand the b, c, and e 3BR from the integer period ratio to the observed period ratio as $t_{\mathrm{exp}}$. In our simulation, $t_{\mathrm{exp}}$ mainly depends on the tidal damping strength on the planets (determined by $Q_{\mathrm{phy}}$) and is hardly affected by other parameters such as the mass ratios of the planets. For example, panels e and f in Figure 6 show a successful expansion with masses of the planets according to mass model M2a in Table 2 and the same $Q_{\mathrm{phy}}$ for all planets. From the orbital expansion phase simulations (see Section 5), the relation between $t_{\mathrm{exp}}$ and $Q_{\mathrm{phy}}$ can be expressed as:
$t_{\mathrm{exp}}\approx 380Q_{\mathrm{phy}}\,\mathrm{Myr},$ (14)
@@
Such results conflict with the formation model of Kepler-221 in two aspects. Firstly, The value of $Q_{\mathrm{phy}}$ is over two orders of magnitudes smaller than the typical value of the tidal quality factor $Q$ for terrestrial planets (Lee et al., 2013; Silburt & Rein, 2015), indicating the existence of other much more efficient dissipation mechanisms to speed up the simulation. Secondly, the reformation process of 3BR requires gentle convergent migration between planets (b, c, e) corresponding to relatively large $Q_{\mathrm{phy}}$ (in Phase III we apply $Q_{\mathrm{phy}}>3$, see Table 3 and Section 4.4). This conflicts with the $Q_{\mathrm{phy}}<2$ constraint derived from the age of the system.
@@
Table 6: Dynamical configuration of the Kepler-402 planetary system according to Rowe et al. (2014).
@@
Goldberg & Batygin (2021) point out that the timescale puzzle could be solved with some additional dissipation mechanism required (for instance, obliquity tides; Millholland & Laughlin 2019). Obliquity tides occur when planets have a large axial tilt (obliquity) (Millholland & Laughlin, 2019) and have been invoked to speed up the orbital expansion in the Kepler-221 system (Goldberg & Batygin, 2021). A non-zero obliquity for planets can be maintained if the planet is in Cassini state where it reaches a secular spin-orbit resonance with synchronized precession of planetary spin and orbital angular momentum. The total energy dissipation rate for a planet with non-zero eccentricity and obliquity can be expressed as (Millholland & Laughlin, 2019):
$\dot{E}(e,\epsilon)=\frac{2K}{1+\cos^{2}\epsilon}\left[\sin^{2}\epsilon+e^{2}\left(7+16\sin^{2}\epsilon\right)\right],$ (15)
@@
Planets are likely to have low obliquity during the disk phase as their angular momentum and spin vectors are aligned with the disk. But during the brief phase of dynamical instability and the resonance reformation afterward, planet b or c could have acquired a mutual inclination with respect to the other planets, which is a requirement for trapping in a Cassini state. Additionally, planet c dominates the tidal expansion process at the same level as planet b due to the large radius of c compared to b (see Equation (6)). Equation (15) demonstrates that even modest obliquity (for example, on planet c, consistent with Section 4.3) can dominate the energy dissipation, and drive the tidal expansion at rates much higher than what can be obtained from tidal damping alone.